A man sold half an egg more than half of his eggs. He then sold half an egg more than half of his remaining eggs. He did the same the third time and then had 7 eggs left. How many eggs did he had at first?
Answer by nyc_function(2741)

If he had x eggs at the start of the 3rd time then
x/2 + 1/2 = 7
x = 13
At the start of the 2nd time he had 15
x/2 + 1/2 = 15
x = 25
At the start of the 1st time he had 31
x/2 + 1/2 = 31
x = 49.